many marine steam propulsion plants achieve a fuel rate of 0.45 pound of fuel per shaft horsepower hour. determine the equivalent fuel rate in kilograms per shaft kilowatt hour
1
Expert's answer
2020-03-23T13:05:04-0400

1 pound is equal to 0.454 kg "\rArr" the rate from the task can be rewritten as "0.45 pounds * 0. 454 \frac{kg}{pound}\approx0.2 kg" per shaft horsepower hour


1 horsepower is equal to 740 watts. So 0.74 kilowatts are equal to 1 horsepower


So, the rate from the task can be rewritten as 0.2 kg per 0.74 shaft kilowatt hour

or

"\frac{0.2kg}{0.74}" per shaft "\frac{0.74kilowatt}{0.74}" hour


"\frac{0.2kg}{0.74}=0.27kg"


so the rate is equal to

"0.27 kg" per shaft kilowatt hour